# references.bib

### v2

@@
(May 2024)
This paper critically evaluates the HESA (Higher Education Statistics Agency) Data Report for the Employer Justified Retirement Age (EJRA) Review Group at the University of Cambridge ([CambridgeHESA2024]), identifying significant methodological flaws and misinterpretations. Our analysis reveals issues such as application of data filters only to the Cambridge sample, inconsistent variable treatment, and erroneous statistical conclusions. The Report suggests EJRA increased job creation rates at Cambridge, but we show Cambridge consistently had lower job creation rates for Established Academic Careers compared to other Russell Group universities, both before and after EJRA implementation in 2011. This suggests that EJRA is not a significant factor driving job creation rates. Since other universities without an EJRA exhibit higher job creation rates, this suggests job creation can be sustained without such a policy. We conclude that the EJRA did not achieve its intended goal of increasing opportunities for young academics and may have exacerbated existing disparities compared to other leading universities. We recommend EJRA be abolished at Cambridge since it does not meet its justified aims and could be viewed as unlawful age discrimination.
@@ 1 Introduction
The University of Cambridge in the United Kingdom employs an Employer Justified Retirement Age (EJRA) policy, mandating retirement for all academic and academic-related staff at the end of the academic year in which they turn 67. Following the abolition of the Default Retirement Age (DRA) by the Coalition government in 2011 following on from the Equality Act (2010), employers were permitted to impose mandatory retirement only on the condition that it could be objectively justified as a proportional means of achieving a legitimate social policy aim. Implemented in 2012, after consultation with its staff, the Cambridge EJRA has four stated goals:
@@
The effectiveness of the EJRA in achieving its aims is predicated upon its effect upon the rate at which vacancies arise from retirement and other causes. Estimates of the effect of the EJRA on the vacancy creation rate (VCR) may be made using models employing simplifying assumptions. A simple model, presented by Paul Ewart during debates in Congregation at Oxford [^1] See https://tinyurl.com/yc68vksc, page 443., assumes a fixed number of posts subject to the EJRA and a uniform age distribution i.e. equal numbers of staff in each year group by age. (The Data Report confirms that this reflects the age distribution also at Cambridge.)
Then, initially assuming that vacancies arise only by retirements at the EJRA when a cohort reaches the retirement age and are filled by younger academics at the average age of appointment, the process can be viewed as a queue where the numbers joining equal the numbers leaving by retirement. If the EJRA was abolished, then the average age at which people retired would increase, at most, by about 3 years, according to survey data. In the steady state, the length of the queue would increase, but the number in each age group would decrease correspondingly to maintain the same number of staff. Hence the number of vacancies each year would decrease by the same amount.
The effect of an EJRA then is to increase the Vacancy Creation rate (VCR) relative to the situation with no mandatory retirement. It is on this basis that the University seeks to justify its EJRA policy as a means to achieving its stated aims. However, since the policy discriminates on the basis of a protected characteristic, age, it is unlawful unless it can be objectively justified as a proportionate means of achieving those aims. Using this simple “queuing” model, it is possible to estimate the effect of the EJRA on the VCR as a measure of its proportionality.
This article critically evaluates the Data Report, identifying significant methodological flaws and misinterpretations. The Data Report suggests that EJRA at the University of Cambridge increased job creation rates. Our analysis reveals issues such as inappropriate data filters, inconsistent variable treatment, and erroneous statistical conclusions. We show that Cambridge consistently had lower job (vacancy) creation rates for Established Academic Careers (EAC) compared to other Russell Group universities, both before and after the EJRA implementation, implying that EJRA is not a significant factor driving job creation rates. Since other Russell Group universities, which do not have an EJRA, exhibit higher job creation rates, this suggests that job creation can be sustained or even thrive without such a policy.
@@ 2 Qualitative assessment of the EJRA
The Ewart Employment Tribunal judged that an increase in vacancy creation rate of between 2 – 4% was “trivial” and not a proportionate means of achieving the aims. [^2] See https://tinyurl.com/2ywaft57 This finding was upheld by the Employment Appeal Tribunal [^3] See https://tinyurl.com/3eea3bkm and was further confirmed in a recent ET cases (2023) involving four claimants at Oxford University. [^4] See https://tinyurl.com/447ecv5a, https://tinyurl.com/5ahknzkk, https://tinyurl.com/5efvjakv, and https://tinyurl.com/3zxt2kd7 It was acknowledged by Oxford University’s statistical analysis in these recent ET tribunals that the modelling method presented by Ewart “gave similar results to the methodology based on Little’s Law of queuing, once the correct assumptions were used.” [^5] Witness statement of Malgorzata Turner in Employment Tribunals referred to in footnote 4 above.
@@
The “steady state” conditions of this simple model do not take account of the transient effect induced by a sudden abolition of the EJRA or a change in the age from say 67 to 69. In the latter case, and using the same assumption of a uniform age distribution, there would be no vacancies for two years after which the situation would evolve into the steady state over time. This effect is treated in a dynamic modelling approach using Little’s Law that would also have to take account of the vacancies created by reasons other than retirement and the average duration of extensions of employment post-EJRA.
It is clear, however, given the similarity in the situations at Oxford and Cambridge; same age at appointment of 40 and 50% of vacancies arising from reasons other than retirement etc., one expects that the increase the vacancy creation rate by the EJRA at Cambridge will also be in the range 2 – 4%. Any significant deviation from this estimate would suggest that some filtering of the statistics data has been made to change the assumptions built into the model.
@@
The EJRA is predicated on an antiquated university model with a fixed number of professorships, necessitating one professor’s departure for another’s ascension. This premise aligns with the “lump of labor fallacy” – the erroneous belief that there is a fixed amount of work to be distributed, leading to the misconception that increased productivity causes unemployment. However, this notion is fallacious ([Krugman2024]). Arguing against productivity enhancements due to potential job losses is irrational. The creation of new professorships is contingent on funding availability and academic interest, not vacancy generation through retirements. Hiring new faculty members can potentially stimulate novel research avenues and ways of thinking, fostering innovation and knowledge creation, rather than merely replacing outgoing scholars. In addition, the senior outgoing scholars are likely able to generate significant amounts of funds, all of which are lost to the university.
@@
It is improbable that junior academics at Cambridge directly benefit from the forced retirement of their senior colleagues. Senior positions are advertised globally, and the potential for internal promotions is often outweighed by external applicants. Thus, even if the model underpinning the Review Report is sound (which it clearly isn’t), the intergenerational fairness argument would not apply within the University context but rather on a global scale, where Cambridge’s retirees ostensibly create opportunities for younger academics worldwide. Needless to say, this would only be a drop in the ocean and by no means a proportionate measure. Curiously, this rationale does not extend to administrative positions, which are exempt from the EJRA. Additionally, the policy disproportionately disadvantages women, whose career trajectories are often prolonged due to family obligations, resulting in a shorter timeframe before reaching the retirement age after achieving career milestones. In figure 1, the university argues that the EJRA policy promoted in the HESA review increases the lectureship opportunities available to post-doctoral colleagues ([PostDocWebsite]).
@@
Since the Data Report suggests increasing the retirement age by two years to 69, it is difficult to understand how lectureship opportunities will be increased. One possibility is that the University plans on increasing the number of senior positions such as professorships. Unfortunately, a simple graph of the number of senior professorships at Cambridge, depicted in figure 2 shows a very sharp decline in these positions after the EJRA period in Cambridge.
@@
In summary, the arguments highlight the potential adverse impacts of the EJRA on innovation, succession planning, intergenerational fairness, and academic freedom, contending that the policy may be counterproductive to the University’s stated aims and could hinder its pursuit of excellence in research and education.
@@ 3 Quantitative assessment of the Data Report
The Data Report sought to establish that EJRA has a beneficial effect on vacancy creation based on historical data and statistical techniques. In general terms and when properly applied, the method of analysis (difference-in-differences) used is capable of measuring the impact of a policy by comparing changes, from before to after the policy change, between a group that experienced the policy and a group that did not. In the analysis of whether the removal of EJRA at Cambridge would affect job creation, one group was the RRG (Remaining Russell Group) where EJRA was removed, and the other group comprised solely of Cambridge where an EJRA was introduced so that mandatory retirement continued, and the difference-in-differences calculation was applied to vacancy creation rates in the RRG and in Cambridge, before and after EJRA removal in the RRG. This section shows that, contrary to sound practice, Cambridge’s data was selectively filtered while other universities’ data were not treated the same way. The analysis as applied did not adequately control for pre-existing trends, or for differences between universities. In fact, Cambridge consistently had lower vacancy creation rates compared to other universities, both before and after EJRA removal in the RRG. It is shown below that the conclusion that the removal of the EJRA policy would significantly impact the vacancy creation rate in Cambridge simply does not follow from the analysis.
@@ 3.1 Effect of EJRA on Job Creation Rate
The authors of the Data Report obtained data from HESA on the number of employees in different categories for each year from 2007-2008 academic year to 2021-2022 academic year for each of the 24 Russell Group universities including Cambridge and Oxford. Oxford is excluded from the analysis with a comparison made only between Cambridge and RRG. The authors describe on page 7 how they applied various “filters” to the HESA data using Cambridge HR data to construct the group that is eligible for EJRA, which was not identified from the HESA data alone. They did not do and could not do such a data editing for the RRG universities and so it is not clear that a fair comparison was made. To characterize this choice, if an RRG university wanted to introduce an EJRA, would they also apply it only to a subset of the individuals recorded in the HESA dataset?
The Data Report examines two classes of staff at Cambridge: Established Academic staff (EAC) and Established Academic-related staff (EAR). The EAC group includes academic staff whose primary function is teaching and/or research. Specifically, the Data Report defines EAC as employees with:
@@
In effect, the EAC group represents the core group of established academic teaching and research staff, while the EAR group covers the non-academic professional support staff, both at relatively senior levels within the university. The purpose of the Data Report is to show that the university created a significantly higher number of jobs in the Cambridge EAC group than would have arisen if Cambridge did not have an EJRA by comparing with other universities who do not operate an EJRA.
@@
Yet another possibility is that something else is to blame - perhaps for example, there were other factors, presumably macroeconomic, that affected job creation over this period. In that case, what we need to do is to compare the average performance of Cambridge post 2011 with the average performance of Cambridge pre-2011 and do the same for the RRG universities. This is the heart of why their argument appears to work since one can see that just prior to 2012 the RRG universities created a lot of new positions and their performance subsequent to 2012 was relatively worse. Actually, the key feature of the pre-2012 data was the big dip following the 2008 Global Financial Crisis (GFC) and the strong rebound, especially for RRG universities, just running up to 2012.
The main approach they use is a so-called natural experiment, which is intended to capture how clinical trials are carried out. The treatment or policy variable is the introduction of the EJRA by Cambridge in 2011, and the treatment group is Cambridge with pre-treatment period 2007/2008-2011/2012 and post treatment period 2012/2013 (it takes a year for the abolition of retirement to take effect) to 2021/2022. The control group is the RRG universities. It is important to note that this design addresses the issue about abolition of EJRA as it compares universities without EJRA with Cambridge that does. It does not provide any evidence about the effect of raising the age of EJRA to 69 from 67. As outlined in the introduction above, increasing the age of the EJRA results in its being even less proportionate as a means of achieving the aims.
The main issues for application of this method here are:
@@
    Treatment is not randomly assigned as in a clinical trial. Cambridge, the authors themselves argue (and the authors of the Oxford EJRA argued before them), is different and those differences lead them to choose the EJRA policy in the first place. Fundamentally, one cannot separate out whether the difference between Cambridge and the RRG universities is due to the EJRA or due to the fundamental difference already present.
@@
    The treatment is really 2012 itself, not the introduction of the EJRA (or rather the division is into the pre- and post-2012 periods). The EJRA was not the only thing that changed after 2012 compared to before 2012 and that could influence vacancy rates.
@@
    The sample sizes are too small. There is only one treated patient, Cambridge, and only 22 untreated patients, the RRG. Imagine testing a vaccine on one patient and applying a placebo to twenty two. Pfizer’s Covid-19 vaccine clinical trial had over 46,000 participants. The statistics here all rely on large sample approximations, and here the smallest group average is made from 5 time series observations.
@@
    The unit of measurement is at the university level. In a clinical trial, one is looking at individuals and it is reasonable to assume that although they are different in fundamental ways, their health outcomes are essentially independent across individuals and only depends on the treatment. It is hard to believe that is the case for university’s economic decisions, which is what is needed here. For example, all the universities are admitting undergraduates through UCAS and they are all trying to recruit from the same pool of academics.
@@
In conclusion, the methods adopted are not appropriate for the data they have. The results they obtained are not credible. The fundamental fact is that Cambridge creates fewer new academic jobs in comparison with RRG whether or not they both had a retirement age. This is clear from Figure A1.
@@
Figure A30 from the Data Report (reproduced in Figure 4) gives their estimates of the treatment effect separately for each year along with confidence bands. This is defined as the difference in job creation between Cambridge and RRG relative to the same difference for 2012 (the zero value for 2012 is suppressed from the figure). According to this graph, the introduction of EJRA had a positive relative effect even five years before it was introduced! In the language of natural experiments, this evidences that the pre treatment parallel trend assumption is violated, which they acknowledge, and which leads them to implement a further couple of inappropriate methods for dealing with this issue. In the Appendix we give a more formal treatment of the statistical issues. Another interesting aspect of this figure is that there is a declining trend in job creation rates in the post-Brexit period (from 2017 onwards) interrupted by a spike upwards during the Covid year of 2020. Even taking their methods at face value, by 2021, the effect has totally disappeared, so whatever was driving the results has vanished. Another interpretation of this data is just that, as in any time series, there are periods where the series is above trend and periods where it is below trend and these cycles are determined by a complex set of reasons that have not been investigated in this report.

### v3

@@
(June 12th, 2024)
This paper critically evaluates the HESA (Higher Education Statistics Agency) Data Report for the Employer Justified Retirement Age (EJRA) Review Group at the University of Cambridge ([CambridgeHESA2024]), identifying significant methodological flaws and misinterpretations. Our analysis reveals issues such as unclear application of data filters, inconsistent variable treatment, and erroneous statistical conclusions. The Report suggests that the EJRA increased job creation rates at Cambridge, but we show Cambridge consistently had lower job creation rates for Established Academic Careers compared to other Russell Group universities, both before and after EJRA implementation in 2011, with no evidence for a significant change in this deficit post implementation. This suggests that EJRA is not a significant factor driving job creation rates. Since other universities without an EJRA exhibit higher job creation rates, this suggests job creation can be sustained without such a policy. We conclude that the EJRA did not achieve its intended goal of increasing opportunities for young academics and may have exacerbated existing disparities compared to other leading universities. We recommend EJRA be abolished at Cambridge since it does not meet its aims and could be viewed as unlawful age discrimination.
This version is a revision reflecting some of the comments made by members of the university EJRA review group in public discussion, see [Holmes24].
@@ 1 Introduction
The University of Cambridge employs an Employer Justified Retirement Age (EJRA) policy, mandating retirement for all academic and academic-related staff at the end of the academic year in which they turn 67. Following the abolition of the Default Retirement Age (DRA) by the UK Coalition government in 2011 following on from the Equality Act (2010), employers were permitted to impose mandatory retirement only on the condition that it could be objectively justified as a proportional means of achieving a legitimate social policy aim. Implemented in 2012, after consultation with its staff, the Cambridge EJRA has four stated goals:
@@
The effectiveness of the EJRA in achieving its aims is predicated upon its effect upon the rate at which vacancies arise from retirement and other causes. Estimates of the effect of the EJRA on the vacancy creation rate (VCR) may be made using models employing simplifying assumptions. A simple model, presented by Paul Ewart during debates in Congregation at Oxford [^1] See https://tinyurl.com/yc68vksc, page 443., assumes a fixed number of posts subject to the EJRA and a uniform age distribution i.e. equal numbers of staff in each year group by age. [^2] The Data Report, [CambridgeHESA2024], confirms that this reflects the age distribution also at Cambridge. Then, initially assuming that vacancies arise only by retirements at the EJRA when a cohort reaches the retirement age and are filled by younger academics at the average age of appointment, the process can be viewed as a queue where the numbers joining equal the numbers leaving by retirement. If the EJRA was abolished, then the average age at which people retired would increase, at most, by about 3 years, according to survey data. In the steady state, the length of the queue would increase, but the number in each age group would decrease correspondingly to maintain the same number of staff. Hence the number of vacancies each year would decrease by the same amount.
The effect of an EJRA then is to increase the Vacancy Creation rate (VCR) relative to the situation with no mandatory retirement. It is on this basis that the University seeks to justify its EJRA policy as a means to achieving its stated aims. However, since the policy discriminates on the basis of a protected characteristic, age, it is unlawful unless it can be objectively justified as a proportionate means of achieving those aims. Using this simple queuing model, it is possible to estimate the effect of the EJRA on the VCR as a measure of its proportionality.
This article critically evaluates the Data Report, [CambridgeHESA2024], identifying significant methodological flaws and misinterpretations. The Data Report suggests that EJRA at the University of Cambridge increased job creation rates. Our analysis reveals issues with inappropriate data filters, inconsistent variable treatment, and erroneous statistical conclusions. In fact, Cambridge consistently had lower job (vacancy) creation rates for Established Academic Careers (EAC) compared to other Russell Group universities, both before and after the EJRA implementation. Furthermore, they supply no credible evidence for a significant change in this deficit after implementation. This implies that EJRA is not a significant factor driving job creation rates. Since other Russell Group universities, which do not have an EJRA, exhibit higher job creation rates, this suggests that job creation can be sustained or even thrive without such a policy.
@@ 2 Qualitative assessment of the EJRA
The Ewart Employment Tribunal judged that an increase in vacancy creation rate of between 2-4% was trivial and not a proportionate means of achieving the aims. [^3] See https://tinyurl.com/2ywaft57 This finding was upheld by the Employment Appeal Tribunal [^4] See https://tinyurl.com/3eea3bkm and was further confirmed in a recent ET cases (2023) involving four claimants at Oxford University. [^5] See https://tinyurl.com/447ecv5a, https://tinyurl.com/5ahknzkk, https://tinyurl.com/5efvjakv, and https://tinyurl.com/3zxt2kd7 It was acknowledged by Oxford University’s statistical analysis in these recent ET tribunals that the modelling method presented by Ewart ”gave similar results to the methodology based on Little’s Law of queuing, once the correct assumptions were used”. [^6] Witness statement of Malgorzata Turner in Employment Tribunals referred to in footnote 5 above.
@@
The ”steady state” conditions of this simple model do not take account of the transient effect induced by a sudden abolition of the EJRA or a change in the age from say 67 to 69. In the latter case, and using the same assumption of a uniform age distribution, there would be no retirement-generated vacancies for two years after which the situation would evolve into the steady state over time. This effect is treated in a dynamic modelling approach using Little’s Law that would also have to take account of the vacancies created by reasons other than retirement and the average duration of extensions of employment post-EJRA.
It is clear, however, given the similarity in the situations at Oxford and Cambridge; same age at appointment of 40 and 50% of vacancies arising from reasons other than retirement etc., one expects that the increase in the vacancy creation rate by the EJRA at Cambridge will also be in the range 2 -4%. Any significant deviation from this estimate would suggest that some filtering of the statistics data has been made to change the assumptions built into the model.
@@
The EJRA is predicated on an antiquated university model with a fixed number of professorships, necessitating one professor’s departure for another’s ascension. This premise aligns with the ”lump of labor fallacy”, the erroneous belief that there is a fixed amount of work to be distributed, leading to the misconception that increased productivity causes unemployment. However, this notion is fallacious ([Krugman2024]). Arguing against productivity enhancements due to potential job losses is irrational. The creation of new professorships is contingent on funding availability and academic interest, not vacancy generation through retirements. Hiring new faculty members can potentially stimulate novel research avenues and ways of thinking, fostering innovation and knowledge creation, rather than merely replacing outgoing scholars. It is ageist prejudice, however, that assumes older academics do not make new or important contributions. Research has shown, for example, that a scientist’s most impactful contribution can come at any stage of his or her career, [Sinatra2016]. In addition, the senior outgoing scholars are likely able to generate significant amounts of funds, all of which are lost to the university.
@@
It is improbable that junior academics at Cambridge directly benefit from the forced retirement of their senior colleagues. Senior positions are advertised globally, and the potential for internal promotions is often outweighed by external applicants. Thus, even if the model underpinning the Review Report is sound (which it clearly isn’t), the intergenerational fairness argument would not apply within the University context but rather on a global scale, where Cambridge’s retirees ostensibly create opportunities for younger academics worldwide. Needless to say, this would only be a drop in the ocean and by no means a proportionate measure. Curiously, this rationale does not extend to administrative positions, which are exempt from the EJRA. Additionally, the policy disproportionately disadvantages women, whose career trajectories are often prolonged due to family obligations, resulting in a shorter time frame before reaching the retirement age after achieving career milestones. In any case, the increase in the number of vacancies each year attributable to the EJRA is trivial - of the order of a few percent
@@
In Figure 1, the university argues that the EJRA policy promoted in the HESA review increases the lectureship opportunities available to post-doctoral colleagues ([PostDocWebsite]). Since the Data Report suggests increasing the retirement age by two years to 69, it is difficult to understand how lectureship opportunities will be increased. One possibility is that the University plans on increasing the number of senior positions such as professorships. Unfortunately, a simple graph of the number of senior professorships at Cambridge, depicted in Figure 2 shows a very sharp decline in these positions after the EJRA period in Cambridge.
@@
In summary, these arguments undermine the claims in the Review Group’s Report, that the EJRA contributes significantly to the aims of intergenerational fairness, innovation, succession planning or academic freedom and, on the contrary, may have adverse effects upon its pursuit of excellence in research and education.
@@ 3 Quantitative assessment of the Data Report
The Data Report sought to establish that EJRA has a beneficial effect on vacancy creation based on historical data and statistical techniques. In general terms and when properly applied, the method of analysis (difference-in-differences) used is capable of measuring the impact of a policy by comparing changes, from before to after the policy change, between a group that experienced the policy and a group that did not. In the analysis of whether the removal of EJRA at Cambridge would affect job creation, one group was the RRG (Remaining Russell Group) where EJRA was removed, and the other group consisted solely of Cambridge where an EJRA was introduced so that mandatory retirement continued, and the difference-in-differences calculation was applied to vacancy creation rates in the RRG and in Cambridge, before and after EJRA removal in the RRG. This section argues that Cambridge’s HR data was apparently used to design the filter that was then applied to the HESA data of all universities, and yet full details of this method were not provided, which is contrary to sound practice. The analysis as applied did not adequately control for pre-existing trends, or for differences between universities. In fact, Cambridge consistently had lower vacancy creation rates compared to other universities, both before and after EJRA removal in the RRG. It is shown below that the conclusion that the removal of the EJRA policy would significantly impact the vacancy creation rate in Cambridge simply does not follow from the analysis.
@@ 3.1 Effect of EJRA on Job Creation Rate
The authors of the Data Report obtained data from HESA on the number of employees in different categories for each year from 2007-2008 academic year to 2021-2022 academic year for each of the 24 Russell Group universities including Cambridge and Oxford. Oxford is excluded from the analysis for political/legal rather than scientific reasons with a comparison made only between Cambridge and RRG. The authors describe on page 7 how they applied various “filters” to the HESA data using Cambridge HR data to construct the group that is eligible for the EJRA (which was not identified from the HESA data alone). Since they do not provide details, it is unclear how they chose the filter to match the Cambridge HESA data with the Cambridge HR data. They then applied that same filter to all of the HESA data for the RRG group as well. They did not and could not obtain HR data from the RRG universities and so again, it is not completely clear that a fair comparison was made given the differences across these diverse universities. [Holmes24] reports that ” the filtered dataset covers 93% of the relevant staff cohorts and excludes all staff not subject to EJRA”. In fact, before 2011 everyone in Cambridge and RRG were subject to the national Default Retirement Age and so it is not so clear why attention is restricted to such a subset of the work force, and whether the exclusionary criteria have the same meaning at different universities. For example, salary point 38 and above is likely to mean something quite different at the LSE versus Exeter University and versus Cambridge. It also matters whether this so-called filter was preregistered or whether it was introduced because the authors’ calculations without the filter were not so impressive.
The Data Report examines two classes of staff at Cambridge: Established Academic staff (EAC) and Established Academic-related staff (EAR). The EAC group includes academic staff whose primary function is teaching and/or research. Specifically, the Data Report defines EAC as employees registered in the HESA dataset with:
@@
In effect, the EAC group represents the core group of established academic teaching and research staff, while the EAR group covers the non-academic professional support staff, both at relatively senior levels within the university. The purpose of the Data Report is to show that the university created a significantly higher number of jobs in the Cambridge EAC group than would have arisen if Cambridge had not had an EJRA by comparing with other universities that did not operate an EJRA over the relevant period.
@@
Yet another possibility is that the EJRA improved Cambridge’s job creation relative to what would have happened if it had likewise eliminated mandatory retirement. In that case, what we need to do is to compare the average performance of Cambridge post 2012 with the average performance of Cambridge pre-2012 and do the same for the RRG universities. This is the heart of why their argument appears to work since one can see that just prior to 2012 the RRG universities created a lot of new positions and their performance subsequent to 2012 was relatively worse. Actually, the key feature of the pre-2012 data was the big dip following the 2008 Global Financial Crisis (GFC) and the strong rebound, especially for RRG universities, just running up to 2012, and for Cambridge starting slower but continuing to 2013.
The main approach they use is a so-called natural experiment, which is intended to mimic how clinical trials are carried out. The treatment or policy variable is the introduction of the EJRA by Cambridge in 2011, and the treatment group is Cambridge with pre-treatment period 2007/2008-2011/2012 and post treatment period 2012/2013 (it takes a year for the abolition of retirement to take effect) to 2021/2022. The control group is the RRG universities. They provide a formal statistical test of whether this treatment effect is zero (there is no effect) versus the alternative that there is an effect. It is important to note that this design addresses the issue about abolition of EJRA as it compares universities without EJRA with Cambridge that does. It does not provide any evidence about the effect of raising the age of EJRA to 69 from 67. As outlined in the introduction above, increasing the age of the EJRA results in its being even less proportionate as a means of achieving the aims.
The main issues for the application of this method here are:
@@
    Treatment is not randomly assigned as in a clinical trial. Cambridge, the authors themselves argue is different, (and the authors of the Oxford EJRA argued the same about Oxford before them) and those differences lead them to choose the EJRA policy in the first place. Fundamentally, one cannot separate out whether the difference between Cambridge and the RRG universities or the change in difference is due to the EJRA or due to the fundamental differences already present that manifest themselves in: the level of the vacancy rate, its variability, and how it responds to external shocks.
@@
    The treatment is really 2012 itself, not the introduction of the EJRA (or rather the division is into the pre- and post-2012 periods). The EJRA was not the only thing that changed after 2012 compared to before 2012 and that could influence vacancy rates differentially across universities.
@@
    The sample sizes are too small. There is only one treated patient, Cambridge, and only 22 untreated patients, the RRG (or equivalently, 22 treated patients and one untreated). Imagine testing a vaccine on one patient and applying a placebo to twenty two. Pfizer’s Covid-19 vaccine clinical trial had over 46,000 participants with thousands in both groups. The so-called statistics here all rely on large sample approximations, and here the smallest group average is made from 5 time series observations.
@@
    The outcome variable is a ratio that is typically a small positive number. One gets the impression from reading statistical textbooks (such as [CoxHink74]) that such a variable should not be approached by linear methods.
  5. 5.
@@
    The unit of measurement is at the university level. In a clinical trial, one is looking at individuals, and it is reasonable to assume that although they are different in fundamental ways, the health outcomes are essentially independent across individuals and only depend on the treatment. It is hard to believe that is the case for a university’s economic decisions, which is what is needed here. For example, all the universities are admitting undergraduates through UCAS and they are all trying to recruit from the same pool of academics. The consequence of this is that there probably is not much statistical gain from looking at the RRG universities separately rather than just focussing on their average performance, i.e., there is really just one patient and one control.
@@
In conclusion, the methods adopted are not appropriate for the data they have. The results they obtained are not credible. The fundamental fact is that Cambridge creates fewer new academic jobs in comparison with RRG whether or not they both had a retirement age. This is clear from Figure A1. There is no convincing evidence that in the counterfactual world where Cambridge had no EJRA that it would generate materially fewer vacancies.
@@
Figure A30 from the Data Report (reproduced in Figure 4) gives their estimates of the treatment effect separately for each year along with confidence bands. This is defined as the difference in job creation between Cambridge and RRG relative to the same difference for 2012 (the zero value for 2012 is suppressed from the figure). According to this graph, the introduction of EJRA had a significant positive relative effect even five years before it was introduced! This can’t be right. In the language of natural experiments, this strongly suggests that the pre treatment parallel trends assumption is violated, and so their main results are not valid. This they acknowledge, which leads them to implement a further couple of methods for dealing with this issue. Specifically, [Holmes24] says ”we prefer the student number adjusted approach” whose results are reported in Table A4, Column 3. Interestingly, these results show a significant positive effect not just for the EAC group but also for the EAR group, so it is surprising that the main report recommends abolishing EJRA for the EAC group, if this is their preferred result. Unfortunately, their cure for non parallel trends is not done correctly and raises further issues. In the Appendix we give a more formal treatment of the statistical issues. Another interesting aspect of Figure 4) is that there is a declining trend in job creation rates in the post-Brexit period (from 2017 onwards) interrupted by a spike upwards during the Covid year of 2020. Even taking their methods at face value, by 2021, the effect has totally disappeared, so whatever was driving their results has vanished. Another interpretation of this data is just that, as in any time series, there are periods where the series is above trend and periods where it is below trend and these cycles are determined by a complex set of reasons that have not been investigated in their report. The authors of the report argue that EJRA is the most likely explanation for the positive treatment effect that they found and that our work does not suggest an alternative explanation. Here, we mention a few possible alternative explanations for the variation of these time series over the fifteen years they examine.
  • The 2008, 2014, and 2021 UK Research Exercises have been argued to lead to hiring booms and busts (where initially many new hires are made thereby temporarily increasing the VCR thereafter increasing the denominator so that in subsequent years the VCR falls).
  • The USS pensions scheme went through various changes in response to projected deficits and tax changes that affected retirement decisions and potentially vacancy creation (in 2011 there was an increase in contributions and a career-average scheme for new members; in 2016 there was an increase in contributions and the final salary section was closed to existing members, with a new defined contribution section for higher salaries; in 2018 the proposal to close career-average scheme leads to industrial action).
  • The big global events like the GFC of 2008, the Brexit referendum of 2016 and its aftermath, the Trump trade war and deterioration in international relations, and the Covid spike of 2020 all occurred during the sample period. They all had major impacts on national and even global macroeconomic quantities so they may also be expected to have had an effect on university vacancy rates. In particular, following the GFC of 2008 many universities in the US and UK had hiring freezes, which lead to hiring opportunities for those that did not have such freezes or for others as those freezes were relaxed.
The key point is that all these factors are likely to have differential effects across universities depending for example on the mix of subjects covered by different universities, their REF strategies, and their exposure to global factors. We note that even if a correlation has been established, which is doubtful, such confounding factors make it impossible to claim that evidence has been found that the EJRA is the cause of any effect. For the EJRA to be lawful, objective evidence must be provided by the employer that the policy achieves the aims in a proportionate way and that the aims could not be achieved by other non-discriminatory measures.